ESSENTIAL COMPONENTS OF AN ALGEBRAIC DIFFERENTIAL EQUATION

EVELYNE HUBERT

Abstract. We present an algorithm to determine the essential singular components of an algebraic differential equation. Geometrically, this corresponds to determining the singular solution that have enveloping properties. The algorithm is practical and efficient because it is factorization free, unlike the previous such algorithm.